The equation for the axis of symmetry written as a decimal
for y=x^2-7x+12 is x= _____ .


Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

The equation for the axis of symmetry written as a decimal
for y=x^2-7x+12 is x= _____ .
------------
It's x = -b/2a
x = 7/2
x = 3.5

Answer by stanbon(75887)

The equation for the axis of symmetry written as a decimal
for y=x^2-7x+12 is x= _____ .
-----------------
You have a quadratic with a = 1 ; b = -7 ; c = 12
Axis of symmetry : x = -b/(2a) = 7/2 = 3.5
=============================================
Cheers,
Stan H.